# Privacy-protected Retrieval-Augmented Generation for Knowledge Graph Question Answering

### v1

@@ Abstract
Large Language Models (LLMs) often suffer from hallucinations and outdated or incomplete knowledge. Retrieval-Augmented Generation (RAG) is proposed to address these issues by integrating external knowledge like that in knowledge graphs (KGs) into LLMs. However, leveraging private KGs in RAG systems poses significant privacy risks due to the black-box nature of LLMs and potential insecure data transmission, especially when using third-party LLM APIs lacking transparency and control. In this paper, we investigate the privacy-protected RAG scenario for the first time, where entities in KGs are anonymous for LLMs, thus preventing them from accessing entity semantics. Due to the loss of semantics of entities, previous RAG systems cannot retrieve question-relevant knowledge from KGs by matching questions with the meaningless identifiers of anonymous entities. To realize an effective RAG system in this scenario, two key challenges must be addressed: (1) How can anonymous entities be converted into retrievable information. (2) How to retrieve question-relevant anonymous entities.
To address these challenges, we propose a novel Abstraction Reasoning on Graph (ARoG) framework including relation-centric abstraction and structure-oriented abstraction strategies. For challenge (1), the first strategy abstracts entities into high-level concepts by dynamically capturing the semantics of their adjacent relations. Hence, it supplements meaningful semantics which can further support the retrieval process. For challenge (2), the second strategy transforms unstructured natural language questions into structured abstract concept paths. These paths can be more effectively aligned with the abstracted concepts in KGs, thereby improving retrieval performance. In addition to guiding LLMs to effectively retrieve knowledge from KGs, these two abstraction strategies also strictly protect privacy from being exposed to LLMs. Experiments on three datasets demonstrate that ARoG achieves strong performance and privacy-robustness, establishing a new practical direction for privacy-protected RAG systems.
@@ 1 Introduction
Large Language Models (LLMs) exhibit remarkable capabilities across a wide range of natural language tasks (Wang et al., 2024). However, their practical deployment remains constrained by critical limitations, including factual inaccuracies (hallucinations) and outdated or incomplete knowledge (Ji et al., 2023). To address these issues, the Retrieval-Augmented Generation (RAG) technique has been proposed to first retrieve factual knowledge from external sources and then incorporate them into LLMs during inference, thereby enhancing the factual grounding and reliability of LLM outputs (Shi et al., 2024).
KGs can store a large amount of factual knowledge. Hence, they are widely used in the RAG system as external knowledge sources to provide supplementary knowledge to LLMs (Sun et al., 2024). The effectiveness of KG-enhanced RAG systems is commonly evaluated through Knowledge Graph Question Answering (KGQA) task. Specifically, given a question, the RAG system aims first to retrieve relevant factual triplets from the KGs and then generate answers based on them.
In practice, many private KGs contain sensitive information like personal data and confidential company details (Chen et al., 2023). When using such private KGs to answer questions requiring privacy details, current RAG systems must retrieve relevant privacy data and expose them to the LLMs. As shown in the Step A of Figure 1, when retrieving triplets to answer the question “Where does Bronny live?”, the factual triplet (Bronny,livesin,L.A.)(Bronny,lives\ in,L.A.) containing sensitive information will be exposed to LLMs. The inherent black-box nature of LLMs and potential insecure data transmission between users and LLM servers highlight the urgent need for robust privacy protections. This need is even greater when using third-party LLM APIs, as users may lack transparency or direct control over how their private data are collected, stored, or processed.
@@
Due to the loss of semantics of entities, previous RAG systems cannot retrieve question-relevant knowledge from KGs by matching questions with the meaningless MIDs. To realize an effective RAG system in this scenario, two key challenges must be addressed: (1) How can anonymous entities be converted into retrievable information. (2) How to retrieve question-relevant anonymous entities.
@@
For challenge (1), the relation-centric abstraction strategy treats the anonymous entities as subject or object nouns and their adjacent relations as predicate verbs, then abstracts the entities into corresponding higher-level concepts. These concepts are generated by LLMs based on the adjacent relations of entities. This strategy captures their in-context semantics and then appending to the MIDs, thereby overcoming challenge (1). For instance, an entity that serves as the subject of relations “time_zones”, “contained_by”, “population” and the object of relation “citytown” can be abstracted into “geographic location”.
@@
To fulfill the complete privacy-protected RAG, our ARoG framework also incorporates Abstraction-driven Retrieval module and Generator module. These two modules ensure robust performance in retrieving question-relevant triplets from KGs and generating answers. To evaluate the effectiveness of our ARoG framework, we conduct experiments in the privacy-protected RAG scenario on three popular yet diverse datasets including WebQSP, CWQ, and GrailQA. Our ARoG framework achieves the state-of-the-art (SoTA) performance on all three datasets.
@@ 2.1 Semantic Parsing for KGQA
Semantic parsing (SP)-based methods treat LLMs as semantic parsers that translate questions into formal queries using labeled exemplars. To address errors in the initial formal queries, Li et al. (2023) utilize entity and relation binders grounded in KGs, Nie et al. (2024) reformulate the generation process into code generation process; and Zhang et al. (2025) adapt generation process to function calling using Condition Graphs (CGs), a variant of KGs. However,these methods do not fundamentally resolve the challenges posed by formatting errors, and the performance is heavily limited by the quality and quantity of labeled exemplars.
2.2 Retrieval-Agumented Generation for KGQA
Compared with SP-based methods, RAG-based methods usually have superior performance. RAG-based methods employ LLMs as retrievers to fetch question-relevant triplets as evidence, and employ LLMs as generators to infer answers with evidence. To improve the quality of evidence, Jiang et al. (2023a) propose an iterative LLM-based “Read-then-Reason” framework; Sun et al. (2024) utilize LLMs to perform beam search on KGs; and Xu et al. (2024) utilize LLMs to supplement additional knowledge. However, previous RAG-based KGQA methods directly feed factual triplets within KGs to LLMs, which may undermine data privacy. The inherent black-box nature of LLMs and potential insecure data transmission between users and LLM servers highlight the urgent need for robust privacy protections. This need is even greater when using third-party LLM APIs, as users may lack transparency or direct control over how their privacy data are collected, stored, or processed.
@@ 3.1 Preliminary
Given an nature language question $q$, the topic entities $E_{q}\ =\ \left\{e_{q}^{1},e_{q}^{2},\cdot\cdot\cdot,e_{q}^{n}\right\}$ (the main entities asked in the question (Qiu et al., 2018)), and a knowledge graph $KG$, the task of KGQA is to predict the answer entities $A_{q}\ =\ \left\{e_{a}^{1},e_{a}^{2},\cdot\cdot\cdot,e_{a}^{n}\right\}$, which can answer the question. Following previous works (Sun et al., 2024), we assume that the topic entities and answer entities are labeled and linked to the corresponding entities in $KG$, i.e., $E_{q},A_{q}\subseteq\mathcal{E}$.
@@
To prevent the LLM from accessing sensitive information within the KG, we anonymize all entities by representing them as MIDs. Specifically, we leverage the Freebase database system (Bollacker et al., 2008) to map each entity to a unique, encrypted MID. These MIDs are meaningless to the LLM. Crucially, to maintain task feasibility, the names of the topic entities explicitly mentioned in the question are available.
@@ 3.2 Overview
(1) Relation-centric Abstraction (RA): To overcome the challenge of How can anonymous entities be converted into retrievable information, we develop the Relation-centric Abstraction module. It utilizes LLMs to transform anonymous entities into abstract concepts and then appends to the anonymous identifies to supplement semantic.
@@ 3.3 Relation-centric Abstraction Module
First, starting from a set of $n$ topic entities, we explore the KG to extract $n$ sets of adjacent relations. For each relation cluster $R_{t}$, we utilize an LLM to identify the $W$ most question-relevant relations, which are denoted as $R_{t,opt}$. The process is expressed as follows.
$R_{\text{t,opt}}=\mathop{\arg\max}_{\begin{subarray}{c}R_{\text{t,opt}}\subseteq R_{t}\\ |R_{\text{t,opt}}|\leq W\end{subarray}}\text{LLM}(R_{t},q,Inst_{rr},E_{rr}),$ (1)
@@
Second, while the entities in the candidate clusters are meaningless, their adjacent relations provide rich in-context semantic information. For a representative entity $e$ in the candidate entity clusters, we treat its adjacent relations as verbs, denoted as $R_{v}$. To filter out irrelevant relations, we employ SentenceTransformer (Reimers and Gurevych, 2020) to select the top-$K$ most relevant relations, denoted as $R_{v,opt}$, based on their cosine similarity to the question $q$. The process is expressed as follows.
$R_{v,opt}=\mathop{\arg\max}_{\begin{subarray}{c}R_{v,opt}\subseteq R_{v}\\ |R_{v,opt}|\leq K\end{subarray}}\sum_{r\in R_{v,opt}}\cos(\text{Emb}(q),\text{Emb}(r)),$ (2)
@@
Since the relations within the KG define the schema structures rather than sensitive information, sharing these relations with LLMs poses minimal privacy risk. Additionally, to minimize LLM calls, we assume that entities within the same entity cluster share a common concept, in line with Jiang et al. (2023b). If an entity is associated with multiple generated abstract concepts, these concepts are combined into a comma-separated sequence.
@@ 3.4 Structure-oriented Abstraction Module
where $Inst_{sa}$ describes the abstraction task in a chain-of-thought (CoT) manner (Wei et al., 2023), $E_{sa}$ denotes multiple exemplars. $Rat$ denotes the rationales that represent the thought process and underpin the generation of $P_{q}$, where both are concurrently produced from a single process.
@@ 3.5 Abstraction-driven Retrieval Module
Specifically, as shown in Part (c) of Figure 2, after the Relation-centric Abstraction module, we gather a set of candidate triplets $T_{q}^{abs}$. Then, we retrieve $W$ most relevant triplets $T_{q,\text{opt}}$ from this set, based on their cosine similarity with the triplets $t_{p}$ in the abstract concept path $P_{q}$. The procedure is outlined as follows.
$T_{q,\text{opt}}^{abs}=\mathop{\arg\max}_{\begin{subarray}{c}T^{abs}_{q,\text{opt}}\subseteq T_{q}^{abs}\\ |T^{abs}_{q,\text{opt}}|\leq W\end{subarray}}\sum_{\begin{subarray}{c}t^{abs}\in T^{abs}_{q,\text{opt}}\\ t_{p}\in P_{q}\end{subarray}}\cos(\text{Emb}(t^{abs}),\text{Emb}(t_{p})).$ (5)
@@ 3.6 Generator Module
The Generator module infers the answers by feeding the question and retrieved relevant triplets into the LLM-based generator, which is similar to that of Sun et al. (2024).
@@ 4.1 Experimental Setup
Datasets We conduct experiments on three KGQA datasets including WebQSP, CWQ and GrailQA. The details of the datasets’ splits and the number of selected samples are provided in Appendix A.1. WebQSP (Yih et al., 2016) contains questions from WebQuestions (Berant et al., 2013) that are answerable by Freebase (Bollacker et al., 2008). CWQ (Talmor and Berant, 2018) extends WebQSP, where most questions require at least 2-hop reasoning. GrailQA (Gu et al., 2021) is a diverse KGQA dataset, where most questions necessitate long-tail knowledge.
@@
Metrics We use exact match accuracy (Hits@1) as evaluation metric, consistent with previous studies (Li et al., 2023; Sun et al., 2024; Chen et al., 2024).
Implementation Details We employ gpt-4o-mini-2024-07-18 as the underlying LLM. For the Structure-oriented Abstraction module and the Generator model, the temperature parameter is set to 0 to ensure deterministic outputs. For the Relation-centric Abstraction module, the temperature is set to 0.4 to introduce variability. Both the frequency penalty and presence penalty are set to 0. The width $W$ and depth $D$ of the retrieval are set to 3, striking a balance between performance and efficiency. We repeat the experiment 3 times and report the average scores. More implementation details are illustrated in the Appendix A.3.
@@ 4.2 Baselines
Pure-LLM methods We select IO (Brown et al., 2020), CoT (Wei et al., 2023) and CoT-SC (Wang et al., 2023) as reference. Because they operate without external KGs and do not tailor for KGQA, they serve as baselines for evaluating the intrinsic reasoning capabilities of LLMs.
SP-based methods We select KB-BINDER (Li et al., 2023) and TrustUQA (Zhang et al., 2025) as representative SP-based methods. For their variants with dynamically exemplars, we claim them as KB-BINDER-R and TrustUQA-R, respectively.
RAG-based methods We select ToG (Sun et al., 2024), PoG (Chen et al., 2024) and GoG (Xu et al., 2024) as representative RAG-based methods. ToG (Sun et al., 2024) uses LLMs to retrieve knowledge triplets via a beam search approach. PoG (Chen et al., 2024) enhances the retrieval process through reflection, memory, and adaptive breadth. In a different vein, GoG (Xu et al., 2024) leverages LLMs as a flexible knowledge source, complementing the KG.

### v2

@@ Abstract
Large Language Models (LLMs) often suffer from hallucinations and outdated or incomplete knowledge. Retrieval-Augmented Generation (RAG) is proposed to address these issues by integrating external knowledge like that in knowledge graphs (KGs) into LLMs. However, leveraging private KGs in RAG systems poses significant privacy risks due to the black-box nature of LLMs and potential insecure data transmission. In this paper, we investigate the privacy-protected RAG scenario for the first time, where entities in KGs are anonymous for LLMs, thus preventing them from accessing entity semantics. Due to the loss of semantics of entities, previous RAG systems cannot retrieve question-relevant knowledge from KGs by matching questions with the meaningless identifiers of anonymous entities. To realize an effective RAG system in this scenario, two key challenges must be addressed: (1) How can anonymous entities be converted into retrievable information? (2) How to retrieve question-relevant anonymous entities?
To address these challenges, we propose a novel Abstraction Reasoning on Graph (ARoG) framework including relation-centric abstraction and structure-oriented abstraction strategies. For challenge (1), the first strategy abstracts entities into high-level concepts by dynamically capturing the semantics of their adjacent relations. Hence, it supplements meaningful semantics which can further support the retrieval process. For challenge (2), the second strategy transforms unstructured natural language questions into structured abstract concept paths. These paths can be more effectively aligned with the abstracted concepts in KGs, thereby improving retrieval performance. In addition to guiding LLMs to effectively retrieve knowledge from KGs, these abstraction strategies also strictly protect privacy from being exposed to LLMs. Experiments on three datasets demonstrate that ARoG achieves strong performance and privacy-robustness, establishing a new practical direction for privacy-protected RAG systems.
Codehttps://github.com/NLPGM/ARoG
@@ 1 Introduction
Large Language Models (LLMs) exhibit remarkable capabilities across a wide range of natural language tasks (Wang_Hu_He_Xu_Liu_Liu_Shen_2024). However, their practical deployment remains constrained by critical limitations, including factual inaccuracies (hallucinations) and outdated or incomplete knowledge (ji_hallucination_survey). To address these issues, the Retrieval-Augmented Generation (RAG) technique has been proposed to first retrieve factual knowledge from external sources and then incorporate them into LLMs during inference, thereby enhancing the factual grounding and reliability of LLM outputs (shi-etal-2024-replug).
KGs can store a large amount of factual knowledge. Hence, they are widely used in the RAG system as external knowledge sources to provide supplementary knowledge to LLMs (sun2024thinkongraph). The effectiveness of KG-enhanced RAG systems is commonly evaluated through Knowledge Graph Question Answering (KGQA) task. Specifically, given a question, the RAG system aims first to retrieve relevant factual triplets from the KGs and then generate answers based on them.
In practice, many private KGs contain sensitive information like personal data and confidential company details (10272537). When using such private KGs to answer questions requiring privacy details, current RAG systems must retrieve relevant privacy data and expose them to the LLMs. As shown in the Step A of Figure 1, when retrieving triplets to answer the question “Where does Bronny live?”, the factual triplet (Bronny,livesin,L.A.)(Bronny,lives\ in,L.A.) containing sensitive information will be exposed to LLMs. The inherent black-box nature of LLMs and potential insecure data transmission between users and LLM servers highlight the urgent need for robust privacy protections. This need is even greater when using third-party LLM APIs, as users may lack transparency or direct control over how their private data are collected, stored, or processed.
@@
Due to the loss of semantics of entities, previous RAG systems cannot retrieve question-relevant knowledge from KGs by matching questions with the meaningless MIDs. To realize an effective RAG system in this scenario, two key challenges must be addressed: (1) How can anonymous entities be converted into retrievable information? (2) How to retrieve question-relevant anonymous entities?
@@
For challenge (1), the relation-centric abstraction strategy treats the anonymous entities as subject or object nouns and their adjacent relations as predicate verbs, then abstracts the entities into corresponding higher-level concepts. These concepts are generated by LLMs based on the adjacent relations of entities. This strategy captures their in-context semantics and then appends to the MIDs, thereby overcoming challenge (1). For instance, an entity that serves as the subject of relations “time_zones”, “contained_by”, “population” and the object of relation “citytown” can be abstracted into “geographic location”.
@@
To fulfill the complete privacy-protected RAG, our ARoG framework also incorporates an Abstraction-driven Retrieval module and a Generator module. These two modules ensure robust performance in retrieving question-relevant triplets from KGs and generating answers. To evaluate the effectiveness of our ARoG framework, we conduct experiments in the privacy-protected RAG scenario on three popular yet diverse datasets including WebQSP, CWQ, and GrailQA. Our ARoG framework achieves the state-of-the-art (SoTA) performance on all three datasets.
@@ 2.1 Semantic Parsing for KGQA
Semantic parsing (SP)-based methods treat LLMs as semantic parsers that translate questions into formal queries using labeled exemplars. To address errors in the initial formal queries, li-etal-2023-shot utilize entity and relation binders grounded in KGs, Nie_Zhang_Wang_Liu_2024 reformulate the generation process into code generation process; and Zhang_Jin_Zhu_Chen_Huang_Wang_Hua_Liang_Chen_2025 adapt generation process to function calling using Condition Graphs (CGs), a variant of KGs. However,these methods do not fundamentally resolve the challenges posed by formatting errors, and the performance is heavily limited by the quality and quantity of labeled exemplars.
2.2 Retrieval-Augmented Generation for KGQA
Compared with SP-based methods, RAG-based methods usually have superior performance. RAG-based methods employ LLMs as retrievers to fetch question-relevant triplets as evidence, and employ LLMs as generators to infer answers with evidence. To improve the quality of evidence, jiang-etal-2023-structgpt propose an iterative LLM-based “Read-then-Reason” framework; sun2024thinkongraph utilize LLMs to perform beam search on KGs; and xu-etal-2024-generate utilize LLMs to supplement additional knowledge. However, previous RAG-based KGQA methods directly feed factual triplets within KGs to LLMs, which may undermine data privacy. The inherent black-box nature of LLMs and potential insecure data transmission between users and LLM servers highlight the urgent need for robust privacy protections. This need is even greater when using third-party LLM APIs, as users may lack transparency or direct control over how their privacy data are collected, stored, or processed.
@@ 3.1 Preliminary
Given an nature language question $q$, the topic entities $E_{q}\ =\ \left\{e_{q}^{1},e_{q}^{2},\cdot\cdot\cdot,e_{q}^{n}\right\}$ (the main entities asked in the question (10.1145/3184558.3186916)), and a knowledge graph $KG$, the task of KGQA is to predict the answer entities $A_{q}\ =\ \left\{e_{a}^{1},e_{a}^{2},\cdot\cdot\cdot,e_{a}^{n}\right\}$, which can answer the question. Following previous works (sun2024thinkongraph), we assume that the topic entities and answer entities are labeled and linked to the corresponding entities in $KG$, i.e., $E_{q},A_{q}\subseteq\mathcal{E}$.
@@
To prevent the LLM from accessing sensitive information within the KG, we anonymize all entities by representing them as MIDs. Specifically, we leverage the Freebase database system (Bollacker-etal-2008-freebase) to map each entity to a unique, encrypted MID. These MIDs are meaningless to the LLM. Crucially, to maintain task feasibility, the names of the topic entities explicitly mentioned in the question are available.
@@ 3.2 Overview
(1) Relation-centric Abstraction (RA): To overcome the challenge of How can anonymous entities be converted into retrievable information, we develop the Relation-centric Abstraction module. It utilizes LLMs to transform anonymous entities into abstract concepts and then appends to the anonymous identifiers to supplement semantic.
@@ 3.3 Relation-centric Abstraction Module
First, starting from a set of $n$ topic entities, we explore the KG to extract $n$ sets of adjacent relations. For each relation cluster $R_{t}$, we prompt an LLM to identify the $W$ most question-relevant relations, which are denoted as $R_{t,opt}$. The process is expressed as follows.
$R_{\text{t,opt}}=\mathop{\arg\max}_{\begin{subarray}{c}S\subseteq R_{t},|S|\leq W\end{subarray}}\text{LLM}(R_{t},q,Inst_{rr},E_{rr}),$ (1)
@@
Second, while the entities in the candidate clusters are meaningless, their adjacent relations provide rich in-context semantic information. For any entity $e$ in the candidate entity clusters, we treat its adjacent relations as verbs, denoted as $R_{v}$. To filter out irrelevant relations, we employ SentenceTransformer (reimers-2020-multilingual-sentence-bert) to select the top-$K$ most relevant relations, denoted as $R_{v,opt}$, based on their cosine similarity to the question $q$. The process is expressed as follows.
$R_{v,opt}=\mathop{\arg\max}_{\begin{subarray}{c}S\subseteq R_{v},|S|\leq K\end{subarray}}\sum_{r\in S}\cos(\text{Emb}(q),\text{Emb}(r)),$ (2)
@@
Since the relations within the KG define the schema structures rather than sensitive information, sharing these relations with LLMs poses minimal privacy risk. Additionally, to minimize LLM calls, we assume that entities within the same entity cluster share a common concept, in line with jiang2023unikgqa. If an entity is associated with multiple generated abstract concepts, these concepts are combined into a comma-separated sequence.
@@ 3.4 Structure-oriented Abstraction Module
where $Inst_{sa}$ describes the abstraction task in a chain-of-thought (CoT) manner (wei2023chainofthoughtpromptingelicitsreasoning), $E_{sa}$ denotes multiple exemplars. $Rat$ denotes the rationales that represent the thought process and underpin the generation of $P_{q}$, where both are concurrently produced from a single process.
@@ 3.5 Abstraction-driven Retrieval Module
Specifically, as shown in Part (c) of Figure 2, after the Relation-centric Abstraction module, we gather a set of candidate triplets $T_{q}^{abs}$. Then, we retrieve $W$ most relevant triplets $T_{q,\text{opt}}^{abs}$ from this set, based on their cosine similarity with the triplets $t_{p}$ in the abstract concept path $P_{q}$. The procedure is outlined as follows.
$T_{q,\text{opt}}^{abs}=\underset{\begin{subarray}{c}S\subseteq T_{q}^{abs},|S|\leq W\end{subarray}}{\arg\max}\;\sum_{t^{abs}\in S}\max_{t_{p}\in P_{q}}\cos(\text{Emb}(t^{abs}),\text{Emb}(t_{p})))$ (5)
@@ 3.6 Generator Module
The Generator module infers the answers by feeding the question and retrieved relevant triplets into the LLM-based generator, which is similar to that of sun2024thinkongraph.
@@ 4.1 Experimental Setup
Datasets We conduct experiments on three KGQA datasets including WebQSP, CWQ and GrailQA. The details of the datasets’ splits and the number of selected samples are represented in Table 1. WebQSP (yih-etal-2016-value) contains questions from WebQuestions (berant-etal-2013-semantic) that are answerable by Freebase (Bollacker-etal-2008-freebase). CWQ (talmor-berant-2018-web) extends WebQSP, where most questions require at least 2-hop reasoning. GrailQA (Gu-etal-2021-grailqa) is a diverse KGQA dataset, where most questions necessitate long-tail knowledge.
Dataset #(Train/Dev/Test) #Total #Filtered
WebQSP 3,098/-/1,639 1,639 535
CWQ 27,639/3,519/3,531 1,000 449
GrailQA 44,337/6,763/13,231 1,000 645
Table 1: Statistics of the datasets.
@@
Metrics We use exact match accuracy (Hits@1) as evaluation metric, consistent with previous studies (li-etal-2023-shot; sun2024thinkongraph; NEURIPS_2024_POG).
Implementation Details We employ gpt-4o-mini-2024-07-18 as the underlying LLM. For the Structure-oriented Abstraction module and the Generator model, the temperature parameter is set to 0 to ensure deterministic outputs. For the Relation-centric Abstraction module, the temperature is set to 0.4 to introduce variability. Both the frequency penalty and presence penalty are set to 0. The width $W$ and depth $D$ of the retrieval are set to 3, striking a balance between performance and efficiency. We repeat the experiment 3 times and report the average scores.
@@ 4.2 Baselines
Pure-LLM methods We select IO (NEURIPS-2020-io), CoT (wei2023chainofthoughtpromptingelicitsreasoning) and CoT-SC (wang2023selfconsistency) as reference. Because they operate without external KGs and do not tailor for KGQA, they serve as baselines for evaluating the intrinsic reasoning capabilities of LLMs.
SP-based methods We select KB-BINDER (li-etal-2023-shot) and TrustUQA (Zhang_Jin_Zhu_Chen_Huang_Wang_Hua_Liang_Chen_2025) as representative SP-based methods. For their variants with dynamically exemplars, we claim them as KB-BINDER-R and TrustUQA-R, respectively.
RAG-based methods We select ToG (sun2024thinkongraph), PoG (NEURIPS_2024_POG) and GoG (xu-etal-2024-generate) as representative RAG-based methods. ToG (sun2024thinkongraph) uses LLMs to retrieve knowledge triplets via a beam search approach. PoG (NEURIPS_2024_POG) enhances the retrieval process through reflection, memory, and adaptive breadth. In a different vein, GoG (xu-etal-2024-generate) leverages LLMs as a flexible knowledge source, complementing the KG.